OpenStudy (anonymous):

y=2x-1 inverse function. show steps PLEASE :D

OpenStudy (anonymous):

y+1 = 2x x = (y+1)/2 then swap x<->y: y = (x+1)/2 is inverse function

OpenStudy (anonymous):

is that showing all the work? @triangleguru

OpenStudy (anonymous):

yes, inverse function means to express x as function of y and then swap x and y in formula. Many functions have conditions on x,y where function and its inverse function exists. Some functions have inverse only on limited intervals (like periodic functions). In this case y = (x+1)/2 exists for any x.